OpenStudy (anonymous):

Hi folks, how can we solve for (a) this kind of Equations : a= -1 mod 7 a= 0 mod 13

ganeshie8 (ganeshie8):

a = 0 mod 13 => a = 13k

OpenStudy (anonymous):

good but why not 7k-1

ganeshie8 (ganeshie8):

plug that value in first equation

OpenStudy (anonymous):

13k=-1 mod 7

ganeshie8 (ganeshie8):

13k = -1 mod 7 14k - k = -1 mod 7 -k = -1 mod 7 k = 1 mod 7

ganeshie8 (ganeshie8):

that gives k = 7m+1

ganeshie8 (ganeshie8):

plug this k value in the earlier 'a' value u got :- a = 13k = 13(7m+1) = 13*7m + 13 = 13 mod (13*7)

ganeshie8 (ganeshie8):

thats the solution to given system of congruences, see if that makes some sense..

OpenStudy (anonymous):

How, you were looking for k and then a ??

OpenStudy (anonymous):

in addition how do you move from this : 14k - k = -1 mod 7 to this : -k = -1 mod 7

ganeshie8 (ganeshie8):

ur ultimate goal is to find a

ganeshie8 (ganeshie8):

ok lets look at one question at a time :) 14k - k = -1 mod 7

OpenStudy (anonymous):

yeah exactly so why you were looking for k

ganeshie8 (ganeshie8):

14k is divisible by 7, so we say 14k = 0 mod 7

ganeshie8 (ganeshie8):

14k - k = -1 mod 7 0 - k = -1 mod 7 k = 1 mod 7

OpenStudy (anonymous):

oh cool is it a rule ?

ganeshie8 (ganeshie8):

oly definition/rule we need to remember is this : \(\large a \equiv b \mod n \) means, \(\large n | (a-b) \)

ganeshie8 (ganeshie8):

\(14k \equiv 0 \mod 7\) so replace 14k with 0, cuz 14k leaves remainder 0 when divided by 7

ganeshie8 (ganeshie8):

for ur first question :- a= 0 mod 13 => 13 | (a-0) 13 | a a = 13k

OpenStudy (anonymous):

i mean changing 14k by 0 how do you define this

ganeshie8 (ganeshie8):

that means, 14k leaves a remainder 0 when divided by 7

ganeshie8 (ganeshie8):

14k = 0 = 28 in mod 7 world

OpenStudy (anonymous):

mmm

OpenStudy (anonymous):

so a = 91k+13

ganeshie8 (ganeshie8):

yes, convince ur self by putting k = 0, 1, 2...

ganeshie8 (ganeshie8):

it should leave 0 when divided by 13, and -1 when divided by 7

OpenStudy (anonymous):

yeah it's Working, in fact i can't even understand a lot of Congruent rules

ganeshie8 (ganeshie8):

takes some time to get use to notation, but there are not many rules related to congruences. reverything boils down to oly one definition : a = b mod n => n | (a-b)

OpenStudy (anonymous):

always i say that the reminder of a/n is b

OpenStudy (anonymous):

so i can replace the number by it's reminder ?

OpenStudy (anonymous):

it's \[\equiv\]

ganeshie8 (ganeshie8):

yes cuz \(a \equiv b \mod n\) => \(n | (a-b) \) that also gives us, \(n | (b-a)\) or \(b \equiv a \mod n \) so u can replace \(a\) and \(b\) in congruences

ganeshie8 (ganeshie8):

in simple words :- if a-b is divisible by n, then b-a is also divisible by n so u can replace a and b

OpenStudy (anonymous):

so: if a=3 mod 7 ab=5 mod 7 i can say that : 3a=5 mod 7 ?

ganeshie8 (ganeshie8):

yes absolutely ! a = b mod n, and c = d mod n, then ac = bd mod n

OpenStudy (anonymous):

yeah but my case there is no (b times d)

ganeshie8 (ganeshie8):

GIven :- a = 3 mod 7 ab = 5 mod 7 a = 7k + 3 (7k+3)b = 5 mod 7 3b = 5 mod 7

ganeshie8 (ganeshie8):

when u have, a = b mod n u can replace \(a\) and \(b\) as u wish in other congruence expressions

OpenStudy (anonymous):

Nice

OpenStudy (anonymous):

but we didn't applied the rule that you have mantioned

ganeshie8 (ganeshie8):

and one thing that may help understand congruences faster :- a = b mod n dont think of it as : a/n leaves a remainder of b

ganeshie8 (ganeshie8):

correct way to think of it is : (a-b) is divisible by n

OpenStudy (anonymous):

this is exactly what i wast thinking of it as

ganeshie8 (ganeshie8):

or a-b = nk

ganeshie8 (ganeshie8):

yeah i see that, that way of thinking will make it hard to understand congruences (personal experience)

OpenStudy (anonymous):

hhh, it's been while looking for somebody who wants to tell me this

OpenStudy (anonymous):

and actually i begin understand whats happening right there

ganeshie8 (ganeshie8):

good :)

OpenStudy (anonymous):

Thank you very much Brother, if you can find any think can help me please Contact me :)

ganeshie8 (ganeshie8):

np :) i see u doing great.. .just it requires more practice ...